OpenStudy (anonymous):

solve for x (3√x-2)-8=8

OpenStudy (anonymous):

\[3\sqrt{x-2}-8=8\]

OpenStudy (anonymous):

please show work

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3\sqrt{x-2}-8=8 }\) \(\normalsize\color{blue}{ 3\sqrt{x-2}-8\color{red}{ +8}=8\color{red}{ +8} }\) \(\normalsize\color{blue}{ 3\sqrt{x-2}=16 }\) \(\normalsize\color{blue}{ 3\sqrt{x-2} \color{darkgoldenrod}{ \div3 } =16\color{darkgoldenrod}{ \div3 } }\) \(\normalsize\color{blue}{\sqrt{x-2}=\Large{ \frac{16}{3} }}\)

OpenStudy (solomonzelman):

let me know if you need more help finishing the prob.

OpenStudy (anonymous):

@SolomonZelman I need help to finish it.

OpenStudy (anonymous):

I did work up to this point but i got stuck right where you where.

OpenStudy (solomonzelman):

square both sides, and you will get \(\normalsize\color{blue}{ x-2=±\frac{\large {256}}{\large {9}} }\) add 2 to both sides \(\normalsize\color{blue}{ x=2±\frac{\large {256}}{\large {9}} }\) you can also do this, ... \(\normalsize\color{blue}{ x=\frac{\large {18}}{\large {9}}±\frac{\large {256}}{\large {9}} }\) \(\normalsize\color{blue}{ x=\frac{\large {18±256}}{\large {9}} }\)